# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ 1. C’7)
:=$\mathcal{H}:=\left\{h_{\theta}:\Omega\to\mathbb{R}:\,\theta\in\Theta\right\}.$
@@
respectively. Let $N\in\mathbb{N}$. For i.i.d. copies $Z_{1},\dots,Z_{N}$ of $X\sim\mathbb{P}^{Z}$ and $\varepsilon_{1},\dots\varepsilon_{N}$ of $\varepsilon\sim\operatorname{N}(0,1)$ independent of $Z$, consider empirical processes either of the form
@@
Similarly to Nickl and Wang [2024], whenever we apply Lemma A.7 in this work with an uncountable set $\Theta$, one can verify that the supremum can be realized as one over a countable subset of it.
Lemma § A.8 (Covering Numbers of Product Spaces).
Let $(\tilde{S},\left\|\cdot\right\|_{\tilde{S}})$ be a normed space. Assume there exists a function $\psi:\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to\mathbb{R}$, such that for any $r,\rho\in\mathbb{R}_{>0}$ the covering number $N(\mathbb{B}_{\tilde{S}}(0,r),\left\|\cdot\right\|_{\tilde{S}},\rho)=\psi(r,\rho)$ is finite. For $K\in\mathbb{N}$ define $S:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\tilde{S}$ equipped with the norm $\left\|s\right\|_{S}:=\left(\sum_{i\in\left\llbracket K\right\rrbracket}\left\|s_{i}\right\|_{\tilde{S}}^{2}\right)^{\frac{1}{2}}$ for any $s=(s_{1},\dots,s_{K})\in S$. We then have
$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\rho)\leq\psi\left(r,\frac{\rho}{\sqrt{K}}\right)^{K}.$
Proof of 11 (Lemma A.8).
By hypothesis, for all $\rho,r\in\mathbb{R}_{>0}$ there exists a covering $\mathcal{C}_{\tilde{S}}:=\{\tilde{s}^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}\subseteq\tilde{S}$ of $\mathbb{B}_{\tilde{S}}(0,r)$ with $N:=\psi(r,\rho)$, i.e. for all $\tilde{s}\in\mathbb{B}_{\tilde{S}}(0,r)$ there exists $i_{\circ}=i_{\circ}(\tilde{s})\in\left\llbracket N\right\rrbracket$ with
$\left\|\tilde{s}-\tilde{s}^{(i_{\circ})}\right\|_{\tilde{S}}\leq\rho.$
Now let $s=(s_{1},\dots,s_{K})\in\mathbb{B}_{S}(0,r)$ be arbitrary chosen, such that
$\left\|s\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}\right\|_{\tilde{S}}^{2}\leq r^{2}.$
Thus, for all $j\in\left\llbracket K\right\rrbracket$, we have $s_{j}\in\mathbb{B}_{\tilde{S}}(0,r)$. Hence, for each $s_{j}$, $j\in\left\llbracket K\right\rrbracket$, there exists $i=i_{j}\in\left\llbracket N\right\rrbracket$, such that $s_{j}\in\mathbb{B}_{\tilde{S}}(\tilde{s}^{(i_{j})},\rho)$. Defining $s^{(i_{1},\dots,i_{K})}:=(\tilde{s}^{(i_{1})},\dots,\tilde{s}^{(i_{K})})\in S$, we have
$\left\|s-s^{(i_{1},\dots,i_{K})}\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}-\tilde{s}^{(i_{j})}\right\|_{\tilde{S}}^{2}<K\rho^{2}.$
@@
$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\sqrt{K}\rho)\leq N^{K}=\psi(r,\rho)^{K}.$
This shows the claim.
In the following lemma we use Lemma A.8 to derive an upper bound for the covering number of Sobolev balls.
Lemma § A.9 (Covering Numbers of Sobolev Spaces).
@@
    $N(\mathcal{F},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha+\kappa}}\right)$
@@
    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O},W)},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha-\eta}}\right)$
    for any $\rho\in\mathbb{R}_{>0}$.
Proof of 12 (Lemma A.9).
  1. i)
    Restricting ourselfes to the case $d_{W}=1$, the claim is already shown in Nickl et al. [2020]. For general $d_{W}\in\mathbb{N}$, the claim follows by an application of Lemma A.8.
  2. i)
    We restrict ourselves to the case $d_{W}=1$. The general claim follows then by Lemma A.8. First, let $\eta=0$, such that $\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})}=\left\|\cdot\right\|_{\infty}$. If $\mathcal{O}$ corresponds to a one-dimensional interval, in chapter 4 of Giné and Nickl [2016], a covering result is shown. Using usual extension arguments, one obtains
    $N(\mathcal{F},\left\|\cdot\right\|_{\infty},\rho)\leq\exp\left(c(\alpha,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha}}\right),$
    which is also used in Nickl [2023]. Now, let $\eta\in\mathbb{R}_{>0}\setminus\mathbb{N}$. Following Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$ is compact, and by Edmunds and Triebel [1996], we further have
    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})},\rho)\leq\exp\left(c\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-\eta}}\right).$
    Now let $\eta=k\in\mathbb{N}$. According to Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is compact, as $\alpha>k+\frac{d}{2}$. Thus, for any $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O})$, we have
    $\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta}(\mathcal{O})}\leq c_{\texttt{em}}\cdot\left\|\theta_{1}-\theta_{2}\right\|_{B_{\infty 1}^{k}(\mathcal{O})}.$
    Thus, it sufficies to find an covering of $H^{\alpha}(\mathcal{O},r)$ w.r.t the $\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})}$-norm. Following the theory in Edmunds and Triebel [1996], we have
    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-k}}\right),$
@@
    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{C^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{c_{\texttt{em}}r}{\rho}\right)^{\frac{d}{\alpha-k}}\right).$
@@
Lemma § A.10 (Properties of regular link functions).
@@
    Let $\theta_{1},\theta_{2}\in C^{0}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\infty}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\infty}.$
  2. i)
    Let $\theta_{1},\theta_{2}\in C^{1}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{C^{1}(\mathcal{O})}\leq c(\Psi)\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{1}(\mathcal{O})}.$
  3. i)
    Let $\theta_{1},\theta_{2}\in\mathbb{L}^{2}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}.$
Proof of 13 (Lemma A.10).
Elementary computations show the claim of all three properties, since by hypothesis $\Psi$ as well as all its derivatives are bounded and hence Lipschitz continuous.

### v3

@@ 4.3 Auxiliary results
$\mathcal{H}:=\left\{h_{\theta}:\Omega\to\mathbb{R}:\,\theta\in\Theta\right\}.$
@@
respectively. Let $N\in\mathbb{N}$. For i.i.d. copies $Z_{1},\dots,Z_{N}$ of $Z\sim\mathbb{P}^{Z}$ and $\varepsilon_{1},\dots\varepsilon_{N}$ of $\varepsilon\sim\operatorname{N}(0,1)$ independent of $Z$, consider empirical processes either of the form
@@
Similarly to Nickl and Wang (2024), whenever we apply lemma 4.7 in this work with an uncountable set $\Theta$, one can verify that the supremum can be realized as one over a countable subset of it.
Lemma § 4.8 (Covering Numbers of Product Spaces).
Let (S~,S~)($\tilde$$\|\cdot\|_{\tilde{S}}$) be a normed space. Assume there exists a function $\psi:\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to\mathbb{R}$, such that for any $r,\rho\in\mathbb{R}_{>0}$ the covering number $N(\mathbb{B}_{\tilde{S}}(0,r),\|\cdot\|_{\tilde{S}},\rho)=\psi(r,\rho)$ is finite. For $K\in\mathbb{N}$ define $S:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\tilde{S}$ equipped with the norm $\|s\|_{S}:=\left(\sum_{i\in\left\llbracket K\right\rrbracket}\|s_{i}\|_{\tilde{S}}^{2}\right)^{\frac{1}{2}}$ for any $s=(s_{1},\dots,s_{K})\in S$. We then have
$N(\mathbb{B}_{S}(0,r),\|\cdot\|_{S},\rho)\leq\psi\left(r,\frac{\rho}{\sqrt{K}}\right)^{K}.$
Proof of lemma 4.8.
By hypothesis, for all $\rho,r\in\mathbb{R}_{>0}$ there exists a covering $\mathcal{C}_{\tilde{S}}:=\{\tilde{s}^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}\subseteq\tilde{S}$ of $\mathbb{B}_{\tilde{S}}(0,r)$ with $N:=\psi(r,\rho)$, i.e. for all $\tilde{s}\in\mathbb{B}_{\tilde{S}}(0,r)$ there exists $i_{\circ}=i_{\circ}(\tilde{s})\in\left\llbracket N\right\rrbracket$ with $\|\tilde{s}-\tilde{s}^{(i_{\circ})}\|_{\tilde{S}}\leq\rho$. Now let $s=(s_{1},\dots,s_{K})\in\mathbb{B}_{S}(0,r)$ be arbitrary chosen, such that
$\|s\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\|s_{j}\|_{\tilde{S}}^{2}\leq r^{2},$
and thus, we have $s_{j}\in\mathbb{B}_{\tilde{S}}(0,r)$ for all $j\in\left\llbracket K\right\rrbracket$. Hence, for each $s_{j}$, $j\in\left\llbracket K\right\rrbracket$, there exists $i=i_{j}\in\left\llbracket N\right\rrbracket$, such that $s_{j}\in\mathbb{B}_{\tilde{S}}(\tilde{s}^{(i_{j})},\rho)$. Defining $s^{(i_{1},\dots,i_{K})}:=(\tilde{s}^{(i_{1})},\dots,\tilde{s}^{(i_{K})})\in S$, we have
$\|s-s^{(i_{1},\dots,i_{K})}\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\|s_{j}-\tilde{s}^{(i_{j})}\|_{\tilde{S}}^{2}<K\rho^{2}.$
@@
$N(\mathbb{B}_{S}(0,r),\|\cdot\|_{S},\sqrt{K}\rho)\leq N^{K}=\psi(r,\rho)^{K}.$
This shows the claim. ∎
In the following lemma we use lemma 4.8 to derive an upper bound for the covering number of Sobolev balls.
Lemma § 4.9 (Covering Numbers of Sobolev Spaces).
@@
    $N(\mathcal{F},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha+\kappa}}\right)$
@@
  2. ii)
    $N(\mathcal{F},\|\cdot\|_{C^{\eta}(\mathcal{O},W)},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha-\eta}}\right)$
    for any $\rho\in\mathbb{R}_{>0}$.
Proof of lemma 4.9.
@@
    Restricting ourselfes to the case $d_{W}=1$, the claim is already shown in the proof of Lemma 19 in Nickl et al. (2020), where a differentiation between the cases $\kappa\geq\frac{1}{2}$ and $\kappa<\frac{1}{2}$ is necessary to find proper $(H^{\kappa}(\mathcal{O}))^{*}$-coverings of $H^{\alpha}_{c}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})$, respectively. For general $d_{W}\in\mathbb{N}$, the claim follows by an application of lemma 4.8.
  2. ii)
    We restrict ourselves to the case $d_{W}=1$. The general claim follows then by lemma 4.8. First, let $\eta=0$, such that $\|\cdot\|_{C^{\eta}(\mathcal{O})}=\|\cdot\|_{\infty}$. For $d=1$, if $\mathcal{O}$ equals $[0,1]$, in Giné and Nickl (2016) a covering result is shown, see Corollary 4.3.38 for instance. Using usual extension arguments, one obtains for the general case
    $N(\mathcal{F},\|\cdot\|_{\infty},\rho)\leq\exp\left(c(\alpha,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha}}\right),$
    which is also used in Nickl (2023). Now, let $\eta\in\mathbb{R}_{>0}\setminus\mathbb{N}$. Following remark 4.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$ is compact, and by Edmunds and Triebel (1996) (Section 3.3) and Triebel (1978) (Section 4.10.3), respectively, we further have
    $N(\mathcal{F},\|\cdot\|_{C^{\eta}(\mathcal{O})},\rho)\leq\exp\left(c\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-\eta}}\right).$
    Now let $\eta=k\in\mathbb{N}$. According to remark 4.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is compact, as $\alpha>k+\frac{d}{2}$. Thus, for any $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O})$, we have
    $\|\theta_{1}-\theta_{2}\|_{C^{\eta}(\mathcal{O})}\leq c_{\texttt{em}}\cdot\|\theta_{1}-\theta_{2}\|_{B_{\infty 1}^{k}(\mathcal{O})}.$
    Thus, it sufficies to find an covering of $H^{\alpha}(\mathcal{O},r)$ w.r.t the $\|\cdot\|_{B_{\infty 1}^{k}(\mathcal{O})}$-norm. Following the theory in Edmunds and Triebel (1996) (Section 3.3) and Triebel (1978) (Section 4.10.3), respectively, we have
    $N(H^{\alpha}(\mathcal{O},r),\|\cdot\|_{B_{\infty 1}^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-k}}\right),$
@@
    $N(H^{\alpha}(\mathcal{O},r),\|\cdot\|_{C^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{c_{\texttt{em}}r}{\rho}\right)^{\frac{d}{\alpha-k}}\right).$
@@
Lemma § 4.10 (Properties of Regular Link Functions).
@@
    Let $\theta_{1},\theta_{2}\in C^{0}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{\infty}\leq c(\Psi)\cdot\|\theta_{1}-\theta_{2}\|_{\infty}.$
  2. ii)
    Let $\theta_{1},\theta_{2}\in C^{1}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{C^{1}(\mathcal{O})}\leq c(\Psi)\cdot(1+\|\theta_{1}\|_{C^{1}(\mathcal{O})}\lor\|\theta_{2}\|_{C^{1}(\mathcal{O})})\cdot\|\theta_{1}-\theta_{2}\|_{C^{1}(\mathcal{O})}.$
  3. iii)
    Let $\theta_{1},\theta_{2}\in\mathbb{L}^{2}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c(\Psi)\cdot\|\theta_{1}-\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}.$
Proof of lemma 4.10.
Elementary computations show all three statements of the Lemma, since by hypothesis $\Psi$ as well as all its derivatives are bounded and hence Lipschitz continuous. ∎
Acknowledgements
The author thanks Richard Nickl and Jan Johannes for their time, patience, and fruitful discussions, as well as for the visit to Richard Nickl’s research group in Cambridge in November 2023, which marked the beginning of this work. This work is funded by Deutsche Forschungsgemeinschaft (DFG, German Re- search Foundation) under Germany’s Excellence Strategy EXC-2181/1-39090098 (the Heidelberg STRUCTURES Cluster of Excellence).